A jet aircraft is traveling at 295 m/s in horizontal flight. The engine takes in air at a
rate of 97.8 kg/s and burns fuel at a rate of
3.75 kg/s. The exhaust gases are ejected at
709 m/s relative to the aircraft.
Find the thrust of the jet engine.
Answer in units of N.
Don't round answer.
1
Expert's answer
2019-12-05T10:19:10-0500

Thrust of the jet engine we find from 2nd Newton law


"F=\frac{\Delta p}{\Delta t}=m_1\Delta v+m_2v_0=m_1(v_0-v)+m_2v_0"


Initial conditions


"m_1=97.8 \frac{kg}{s}; m_2=3.75 \frac{kg}{s}; v_0=709\frac{m}{s};v=295\frac{m}{s}"

Calculating

"F=97.8\cdot (709-295)+3.75 \cdot 709= 43148 (N)"